OpenStudy (anonymous):

How do you find the determinant of this matrix by hand? 3, 2, 1 1, 0, -1 0, 1, 2

OpenStudy (amistre64):

3*[..] - 2*[..]+1*[..]

OpenStudy (amistre64):

ive seen one way that takes the first 2 columns and copies them to the right side; then does a diagonal

OpenStudy (anonymous):

do you multiply or do add them?

OpenStudy (amistre64):

3 2 1 3 2 1 0 -1 1 0 0 1 2 0 1 --------------------- 3 2 1 0 -1 1 2 0 1 --------------------- 0+0+1 = 1 1 3 2 0 -1 1 0 1 2 --------------------- 0-3+4 = 1

OpenStudy (amistre64):

the left side I believe is the negative so we get: 1/-1 = -1

OpenStudy (amistre64):

to compare this with a more normal method ....

OpenStudy (amistre64):

3 0 -1 1 2 -------- 3(0(2)-1(-1)) = 3(1) = 3 2 1 -1 0 2 ---------- -2(1(2)-0(-1) = -2(2) = -4 1 1 0 0 1 -------- 1(1(0)-0(1)) = 1(0) = 0 add up the results: 3-4+0 = -1